OpenStudy (anonymous):

HELP ME PLEASE: If tan θ =(√3)/3, 0◦< θ < 360◦, find the larger θ in degrees. and radians

OpenStudy (anonymous):

tan theta will be positive in 3rd quadrant arctan (√3)/3 = 30 degrees larger angle = 180 + 30 = 210 degrees

OpenStudy (anonymous):

what about in radians?

OpenStudy (anonymous):

oh ok 180 degrees = pi radians so 210degrees = 210 pi / 180 = 7pi / 6 radiians